Prove that Re(z)=(z+z')/2
Found 2 solutions by rothauserc, Edwin McCravy:
Answer by rothauserc(4718)

z is the complex number, z = x + yi
:
z' = Re(z) - i*Im(z) where Re(z) is the real portion of z and Im(z) is imaginary portion of z and i is square root of (-1)
:
z + z' = x + yi + Re(z) - i*Im(z)
:
z + z' = x + yi + x - i*y
:
z + z' = x + yi + x - yi = 2x
:
******************
2x / 2 = x = Re(z)
******************
:

Answer by Edwin McCravy(20054)